# ThinK: Thinner Key Cache by Query-Driven Pruning

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications by leveraging increased model sizes and sequence lengths. However, the associated rise in computational and memory costs poses significant challenges, particularly in managing long sequences due to the quadratic complexity of the transformer attention mechanism. This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference. Unlike existing approaches that optimize the memory based on the sequence lengths, we uncover that the channel dimension of the KV cache exhibits significant redundancy, characterized by unbalanced magnitude distribution and low-rank structure in attention weights. Based on these observations, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels. Our approach not only maintains or enhances model accuracy but also achieves a reduction in memory costs by over 20% compared with vanilla KV cache eviction methods. Extensive evaluations on the LLaMA3 and Mistral models across various long-sequence datasets confirm the efficacy of ThinK, setting a new precedent for efficient LLM deployment without compromising performance. We also outline the potential of extending our method to value cache pruning, demonstrating ThinK’s versatility and broad applicability in reducing both memory and computational overheads.
@@ 1 Introduction
Large language models (LLMs) (Hadi et al.,, 2023; Brown et al.,, 2020; OpenAI,, 2023; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Scao et al.,, 2022; Reid et al.,, 2024) have emerged as a dominant paradigm in natural language processing, achieving state-of-the-art performance across various tasks. A key principle, the Scaling Law (Kaplan et al.,, 2020), suggests that LLMs exhibit emergent abilities as model size increases, enhancing their capacity to understand context and handle long sequences (Xiong et al.,, 2023). This capacity growth allows LLMs to generate coherent and contextually accurate responses and enables various downstream applications, such as document summarization (Zhang et al.,, 2019; Zhang et al., 2024a, ), code generation (Chen et al., 2021b, ), and conversational AI (Bordes et al.,, 2016; OpenAI,, 2022), .
@@
Despite their success in various applications, the generation of LLMs incurs significant expenses, which escalate with increasing model size and sequence length. Notably, both the training (Strubell et al.,, 2020; Hoffmann et al.,, 2022; Dong et al., 2024a, ) and inference (Ainslie et al.,, 2023) stages involve frequent generation by LLMs, further contributing to these costs. Consequently, efficient LLMs have gained popularity in recent years (Hu et al.,, 2021; Wan et al.,, 2023). To address these challenges, quantization (Frantar et al.,, 2022; Lin et al.,, 2024; Dettmers et al.,, 2024; Xu et al.,, 2023) and pruning methods (Frankle and Carbin,, 2018; Blalock et al.,, 2020) are employed to reduce model size. Additionally, managing long sequences presents another cost due to the transformer attention mechanism. The quadratic complexity of the attention mechanism results in substantial computational burden when dealing with long sequences, scaling poorly with sequence length. Therefore, effective management of long sequences is essential for the practical deployment of LLMs. In this paper, we focus on the long-context scenario and aim to reduce the memory consumption associated with lengthy sequences.
@@
Specifically, the number of KV cache parameters is the product of batch size $B$, sequence length $S$, number of layers $L$, number of heads $N$, channel size of each head $D$, i.e., $\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times S\times L\times N\times D}$, which need to be stored in the GPU memory during inference. To reduce memory and computational costs during inference, efficiency can only be achieved by pruning the dimensions across $S,L,N,D$ or performing quantization over the caches. It is well-acknowledged that token importance tends to be sparse. Consequently, KV eviction algorithms have been proposed to minimize the KV cache memory cost from dimension $S$ (Xiao et al., 2023b, ; Li et al.,, 2024; Zhang et al., 2024c, ; Leviathan et al.,, 2023). Additionally, inter-layer redundancy has been explored (Liu et al., 2024a, ; Wu and Tu,, 2024; Brandon et al.,, 2024) to address the layer dimension $L$. Despite these advances, existing methods have largely overlooked the channel dimension $D$. In this paper, we highlight that the magnitude across Key cache channel dimensions is significantly unbalanced. Additionally, we observe that the attention weights exhibit a low-rank structure. Based on these findings, we hypothesize that the channel dimension of the key cache exhibits redundancy. Consequently, we explore the redundancy of KV cache tensors specifically from dimension $D$, aiming to devise effective strategies that reduce costs without compromising performance.
@@
In our paper, we propose a simple yet effective method – ThinK, for KV cache pruning. To pinpoint the least significant channels, we formulate the task as an optimization problem, aiming to minimize the loss in attention weights attributable to pruning. To effectively address this problem, we establish a novel query-dependent criterion that assesses the importance of each channel. Using this criterion, we then select the most critical channels in a greedy fashion. We evaluate ThinK using the LLaMA3 (Meta,, 2024) and Mistral (Jiang et al.,, 2023) models and validate its effectiveness across various long-sequence datasets. The results indicate that when paired with token eviction methods, ThinK not only achieves comparable or superior accuracy but also reduces KV cache memory costs by more than 20%.
Contributions.
@@
This work pioneers the investigation into the sparsity structure of the channels in the KV cache. Specifically, we discovered that the activated key cache is sparse given a specific query. This insight allows us to prune the key cache channels using a query-induced norm. Inspired by this phenomenon, we introduce a plug-and-play method ThinK, the first channel pruning method specifically designed for KV cache. ThinK reduces the dimensionality of the cache channel, leading to linear savings in both memory and computational requirements. As plug-and-play techniques, our proposed ThinK is orthogonal to other KV cache compression schemes (e.g. KV cache eviction, quantization). Our extensive experiments demonstrate the remarkable efficiency of ThinK on LLaMA-3 and Mistral models. Furthermore, we explore the promising potential of extending ThinK to value cache pruning (ThinKV) in specific scenarios, highlighting the broad applicability of our method.
@@
Figure 1: An illustration of the pruning procedure of ThinK. Within each head, scores are calculated for each channel, and only the top T channels out of D are selected for retention. A binary channel mask, along with the pruned keys, is subsequently stored in the cache memory.
@@ 2 Observations
We identify several key observations that motivate our approach to prune the channel of KV-Cache. Specifically, we visualize the magnitude of KV-cache and perform singular value decomposition of the attention in LLaMA3-8B model. We show the visualization for certain heads and layers for illustration purposes.
Magnitude of KV-Cache Channel
@@
Figure 2 visualizes the absolute values of key and value cache across the tokens in each channel [^1] We use the visualization code from https://github.com/jy-yuan/KIVI/tree/main/vis.. Consistent with previous findings (Lin et al.,, 2024; Xiao et al., 2023a, ; Liu et al., 2024b, ), we observe that only certain channels have significant magnitudes in the key cache, whereas the value cache lacks obvious patterns. For instance, the absolute value around $50^{th}$ channel of key cache has much larger magnitudes than other channels for all tokens in layer $14$ (Figure 2 (a)). The same observation applies to $50^{th}$ and $150^{th}$ channels of the first head in layer $20$ (Figure 2 (c)). Given such an observation, Liu et al., 2024b proposed to perform quantization over the channels of the key cache based on such observations. In addition to channel quantization, the observation also suggests that it is feasible to prune away certain key cache channels that contribute less to the attention mechanism. Moreover, channel quantization and pruning techniques are orthogonal such that they can be utilized concurrently to enhance model efficiency.
(a) Key Cache ($14^{th}$ layer)
(b) Value Cache ($14^{th}$ layer)
(c) Key Cache ($20^{th}$ layer)
(d) Value Cache ($20^{th}$ layer)
Figure 2: Magnitude of key and value cache for LLaMA3-8B. The first head of layer 14 and layer 20 of LLaMA3-8B is selected to visualize the magnitude of the key and value caches. We observe that the magnitudes of the key cache channels vary differently, whereas the channels of the value cache do not exhibit such variation.
Singular Value Analysis
@@
We conducted singular value decomposition (SVD) (Demmel,, 1997) on the attention of the specified layers to investigate their principal components. The singular values derived from SVD capture the effective rank of the attention matrix, indicating how the information is distributed across different components.
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}$ $=\text{SVD}\bigg{(}\text{softmax}\Big{(}\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{D}}\Big{)}\bigg{)},\quad\text{Energy}_{i}=\frac{\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}$ (1)
@@
Figure 3 (a) illustrates the energy distribution of the singular values, plotted in log scale to enhance the visibility of differences among them. Notably, only a few singular values exhibit high energy levels exceeding $0.01$ across all heads and layers, highlighting their relative significance. This observation aligns with previous findings (Bhojanapalli et al.,, 2021), where a small subset of singular values often captures most of the information in attention mechanisms. In addition, the rapid decay of the energy suggests that a low-rank approximation can effectively capture the essential information in the key cache.
@@
Figure 3 (b), the normalized cumulative energy sum reveals that the top $50$ singular values account for over $90$% of the total energy. These findings suggest that the attention matrix is inherently low-rank (Wang et al.,, 2020; Chen et al., 2021a, ; Dong et al., 2024b, ), indicating that the key cache can be approximated using low-dimensional vectors (Singhania et al.,, 2024).
(a) Energy of singular values
(b) Cumulative energy of singular values
Figure 3: The energy and cumulative energy of the singular values.
@@ 3.1 Preliminary Study of KV Cache Optimization
In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size. Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows:
$2\times B\times S\times L\times N\times D,$ (2)
where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension. The KV cach size grows linearly as the batch size $B$ and sequence length $S$. For a model with multihead attention (MHA) (Vaswani et al.,, 2017), such as LLaMA2-7B (Touvron et al., 2023b, ), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters. The KV cache must be transferred from off-chip memory (HBM) (Jia et al.,, 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck. This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies. Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b, ; Zhang et al., 2024c, ; Li et al.,, 2024) and precision (Hooper et al.,, 2024; Liu et al., 2024b, ). We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
Table 1: Performance comparison of pruning K cache by $l_{p}$ norm on LongBench.
@@
H2O 0 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+$l_{1}$ 30 23.38 17.15 34.99 40.56 31.49 19.90 21.37 22.13 23.44 40.50 90.10 40.65 5.41 69.00 58.64 54.99 37.11
+$l_{1}$ 40 23.51 15.40 34.37 40.71 31.28 20.24 21.25 22.29 22.54 38.50 89.22 39.27 5.87 68.33 58.47 54.33 36.60
+$l_{2}$ 30 23.98 17.04 35.19 39.27 31.29 20.4 21.62 22.46 23.34 40.5 89.75 40.71 5.54 68.67 60.12 58.52 37.40
+$l_{2}$ 40 23.76 16.23 32.19 40.23 32.13 20.69 21.3 22.25 23.2 39.5 89.61 40.24 5.66 69.0 60.09 59.45 37.22
SnapKV 0 24.84 23.96 38.77 42.75 34.55 20.87 22.26 22.61 23.97 70.00 90.52 40.29 5.81 69.50 59.04 51.81 40.10
+$l_{1}$ 30 24.43 24.63 40.11 41.83 33.47 21.22 21.47 22.41 23.73 66.50 90.39 40.20 5.70 68.10 61.04 55.37 40.04
+$l_{1}$ 40 24.58 24.87 39.30 42.76 31.95 20.47 20.95 22.22 23.42 55.50 90.22 39.13 5.82 68.39 60.71 56.10 39.15
+$l_{2}$ 30 24.47 24.73 38.16 41.86 32.23 20.23 21.59 22.45 23.77 67.5 90.33 40.31 5.7 68.42 62.65 60.07 40.28
+$l_{2}$ 40 24.52 23.75 38.35 42.42 32.96 20.39 21.21 22.28 23.41 60.00 90.20 39.59 5.75 68.29 61.96 60.59 39.74
Magnitude base Pruning: Based on the observations in Figure 2 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
@@
Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head:
$I=\textbf{Top}_{T}(M,T)$ (4)
where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel. $n\in[1,N]$ and $d\in[1,D]$ are indicators of heads and channels in key cache. $I\in\mathbb{(Z^{+})}^{N\times T}$ stores the indicators of the top $T$ values in tensor $M$ per head.
In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA3-8B model. We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al.,, 2023). Compared to the baseline methods, H2O (Zhang et al., 2024c, ) and SnapKV (Li et al.,, 2024), both with a KV length of 512, we further prune the channels of the key cache. A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning. The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension. These results also indicate the need for a better pruning matrix to achieve higher pruning ratios effectively.

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications. However, their increased computational and memory demands present significant challenges, especially when handling long sequences. This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference. Unlike existing approaches that optimize the memory based on the sequence length, we identify substantial redundancy in the channel dimension of the KV cache, as indicated by an uneven magnitude distribution and a low-rank structure in the attention weights. In response, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels. Our approach not only maintains or enhances model accuracy but also achieves a reduction in KV cache memory costs by over 20% compared with vanilla KV cache eviction and quantization methods. For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU. Extensive evaluations on the LLaMA and Mistral models across various long-sequence datasets verified the efficiency of ThinK, establishing a new baseline algorithm for efficient LLM deployment without compromising performance.
@@ 1 Introduction
Large language models (LLMs) (Hadi et al., 2023; Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a; b; Scao et al., 2022; Reid et al., 2024) have emerged as a dominant paradigm in natural language processing, achieving state-of-the-art performance across various tasks. A key principle, the Scaling Law (Kaplan et al., 2020), suggests that LLMs exhibit emergent abilities as model size increases, improving their capacity to understand complex context and handle long sequences (Xiong et al., 2023). This growth in capacity enables LLMs to generate coherent, contextually accurate responses and supports a variety of downstream applications, such as document summarization (Zhang et al., 2019; 2024a), code generation (Chen et al., 2021b), solving mathematical problems (Hendrycks et al., 2021; Zhou et al., 2023; Wang et al., 2023; Lightman et al., 2023), and conversational AI (OpenAI, 2022; 2023).
@@
Despite their success in various applications, generating outputs with LLMs incurs significant computational and financial costs, which rise with increasing model size and sequence length. Both the training (Strubell et al., 2020; Hoffmann et al., 2022; Dong et al., 2024a) and inference (Ainslie et al., 2023) stages involve frequent generation, further contributing to these costs. Consequently, efficient LLMs have gained traction in recent years (Hu et al., 2021; Wan et al., 2023). To address these challenges, quantization (Frantar et al., 2022; Lin et al., 2024; Dettmers et al., 2024; Xu et al., 2023) and pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023) are employed to reduce model size. Additionally, the key-value (KV) cache, stored in GPU memory alongside model parameters, scales linearly with both sequence length and batch size, creating a substantial memory burden when handling long sequences. Consequently, effective management of extended contexts is essential for the practical deployment of LLMs. In this paper, we focus on the long-context scenario, aiming to reduce memory consumption associated with processing lengthy sequences.
@@
Specifically, the number of KV cache parameters is the product of batch size $B$, sequence length $S$, number of layers $L$, number of heads $N$, channel size per head $D$, i.e., $\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times S\times L\times N\times D}$, which need to be stored in the GPU memory during inference. To reduce memory and computational costs during inference, efficiency can only be achieved by pruning the dimensions across $S,L,N,D$ or applying quantization to the caches. It is well-acknowledged that token importance tends to be sparse. Consequently, KV eviction algorithms have been proposed to reduce the memory footprint by addressing the sequence length dimension $S$ (Xiao et al., 2023b; Li et al., 2024; Zhang et al., 2024c; Leviathan et al., 2023). Additionally, inter-layer redundancy has been explored (Liu et al., 2024a; Wu & Tu, 2024; Brandon et al., 2024) to address the layer dimension $L$. Despite these advances, existing methods have largely overlooked the channel dimension $D$. In this paper, we highlight that the magnitudes across key cache channel dimensions are significantly imbalanced, and we observe a low-rank structure in attention weights. Based on these findings, we hypothesize that the channel dimension of the key cache exhibits redundancy. Consequently, we focus on exploring the redundancy in the KV cache along dimension $D$, aiming to develop strategies that reduce memory costs without compromising performance.
@@
In this paper, we introduce ThinK, a simple yet effective method for KV cache pruning. To pinpoint the least significant channels, we formulate the problem as an optimization task, aiming to minimize the loss in attention weights caused by pruning. To effectively address this problem, we propose a novel query-dependent criterion that assesses the importance of each channel. Using this criterion, we then select the most critical channels in a greedy fashion. We evaluate ThinK using the LLaMA (Meta, 2024) and Mistral (Jiang et al., 2023) models, and validate its effectiveness across various long-sequence datasets. The results indicate that, when paired with token eviction and KV cache quantization methods, ThinK not only maintains comparable or superior accuracy but also reduces KV cache memory costs by over 20%.
@@
Contributions. This work pioneers the investigation into the sparsity within the channels of the KV cache. Specifically, we uncover that the activated key cache is sparse for a given query. This insight allows us to prune the key cache channels using a query-induced norm. Building on this insight, we introduce ThinK, the first channel pruning method specifically designed for KV cache. ThinK reduces the dimensionality of the cache channels, leading to linear savings in memory usage. As a plug-and-play technique, ThinK is orthogonal to other KV cache compression schemes (e.g. KV cache eviction, quantization). Our extensive experiments demonstrate ThinK’s remarkable efficiency on the LLaMA and Mistral models. Moreover, we explore the potential extension of ThinK to value cache pruning (ThinKV), showcasing the broad applicability of our method.
@@
Figure 1: An illustration of the pruning procedure of ThinK. Within each attention head, scores are computed for each channel, and only the top $T$ channels out of $D$ are selected for retention. A binary channel mask, along with the pruned keys, is then stored in the cache memory.
@@ 2 Observations
We identify several key observations that motivate our approach to pruning the channels of the KV cache. Specifically, we visualize the magnitude of the KV cache and perform singular value decomposition (SVD) on the attention mechanism of the LLaMA model.
@@
Magnitudes of KV cache channels. Figure 4 (in Appendix A) visualizes the absolute values of the KV cache across tokens in each channel [^1] We use the visualization code from https://github.com/jy-yuan/KIVI/tree/main/vis.. Consistent with previous findings (Lin et al., 2024; Xiao et al., 2023a; Liu et al., 2024b), we observe that only certain channels have significant magnitudes in the key cache, whereas the value cache lacks obvious patterns. For instance, in layer $14$ (Figure 4 (a)), the magnitudes in the key cache are substantially higher around the $50^{th}$ channel across all tokens. A similar pattern is observed in the $50^{th}$ and $150^{th}$ channels of the first head in layer $20$ (Figure 4 (c)). Given such an observation, Liu et al. (2024b) proposed to perform quantization over the channels of the key cache. Beyond quantization, our findings suggest that certain key cache channels with smaller contributions to the attention mechanism can be pruned. Moreover, channel quantization and pruning are orthogonal techniques, meaning they can be applied concurrently to further improve model efficiency.
@@
Singular value analysis. We conducted singular value decomposition (SVD) (Demmel, 1997) on the attention weights of the specified layers to investigate their principal components. The singular values derived from SVD capture the effective rank of the attention matrix, indicating how the information is distributed across different components.
@@
Figure 5 (a) (in Appendix A) illustrates the energy distribution of the singular values, plotted on a logarithmic scale to enhance visibility of the differences. Notably, only a few singular values exhibit high energy levels exceeding $0.01$ across all heads and layers, highlighting their relative significance. This observation aligns with previous findings (Bhojanapalli et al., 2021), where a small subset of singular values often captures most of the information in attention mechanisms. In addition, the rapid decay of the energy suggests that a low-rank approximation can effectively capture the essential information in the key cache.
@@
Figure 5 (b) (in Appendix A), the normalized cumulative energy sum reveals that the top $50$ singular values account for over $90$% of the total energy. These findings suggest that the attention matrix is inherently low-rank (Wang et al., 2020; Chen et al., 2021a; Dong et al., 2024b), indicating that the key cache can be approximated using low-dimensional vectors (Singhania et al., 2024).
@@ 3.1 Preliminary Study of KV Cache Optimization
In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size. Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows: $2\times B\times S\times L\times N\times D$, where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension. The KV cach size grows linearly as the batch size $B$ and sequence length $S$. For a model with multihead attention (MHA) (Vaswani et al., 2017), such as LLaMA2-7B (Touvron et al., 2023b), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters. The KV cache must be transferred from off-chip memory (HBM) (Jia et al., 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck. This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies. Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b; Zhang et al., 2024c; Li et al., 2024) and precision (Hooper et al., 2024; Liu et al., 2024b). We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
Table 1: Performance comparison of pruning key cache by $l_{p}$ norm on LongBench.
@@
H2O 0.0 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+$l_{1}$ 0.3 23.38 17.15 34.99 40.56 31.49 19.90 21.37 22.13 23.44 40.50 90.10 40.65 5.41 69.00 58.64 54.99 37.11
+$l_{1}$ 0.4 23.51 15.40 34.37 40.71 31.28 20.24 21.25 22.29 22.54 38.50 89.22 39.27 5.87 68.33 58.47 54.33 36.60
+$l_{2}$ 0.3 23.98 17.04 35.19 39.27 31.29 20.40 21.62 22.46 23.34 40.50 89.75 40.71 5.54 68.67 60.12 58.52 37.40
+$l_{2}$ 0.4 23.76 16.23 32.19 40.23 32.13 20.69 21.30 22.25 23.20 39.50 89.61 40.24 5.66 69.00 60.09 59.45 37.22
SnapKV 0.0 24.84 23.96 38.77 42.75 34.55 20.87 22.26 22.61 23.97 70.00 90.52 40.29 5.81 69.50 59.04 51.81 40.10
+$l_{1}$ 0.3 24.43 24.63 40.11 41.83 33.47 21.22 21.47 22.41 23.73 66.50 90.39 40.20 5.70 68.10 61.04 55.37 40.04
+$l_{1}$ 0.4 24.58 24.87 39.30 42.76 31.95 20.47 20.95 22.22 23.42 55.50 90.22 39.13 5.82 68.39 60.71 56.10 39.15
+$l_{2}$ 0.3 24.47 24.73 38.16 41.86 32.23 20.23 21.59 22.45 23.77 67.50 90.33 40.31 5.70 68.42 62.65 60.07 40.28
+$l_{2}$ 0.4 24.52 23.75 38.35 42.42 32.96 20.39 21.21 22.28 23.41 60.00 90.20 39.59 5.75 68.29 61.96 60.59 39.74
Magnitude based Pruning: Based on the observations in Figure 4 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
@@
Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head: $I=\textbf{Top}_{T}(M,T)$ where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel. $n\in[1,N]$ and $d\in[1,D]$ are indicators of heads and channels in key cache. $I\in\mathbb{(Z^{+})}^{N\times T}$ stores the indicators of the top $T$ values in tensor $M$ per head.
In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA-3-8B model. We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al., 2023). Compared to the baseline methods, H2O (Zhang et al., 2024c) and SnapKV (Li et al., 2024), both with a KV length of 512, we further prune the channels of the key cache. A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning. The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension. These results also indicate the need for a better pruning matrix to achieve higher pruning ratios effectively.